OpenStudy (anonymous):

cynthia_franco : 5x-2= 2(3x-4) 5x-2= 6x-8 -5x -5x -2= x- 8 +8 +8 10 = x

OpenStudy (anonymous):

whoops x= 6

OpenStudy (anonymous):

can double check your answer by plugging 6 into original equation... 5(6) - 2 = 2(3(6)-4) 30- 2 = 2(14) 28= 28